Change page style: 

Gemini-S still recovering from last week's storm

Further to the previous note about the bad weather, because of persistent problems with the electrical power on site it was decided on Aug 13 to shut down all system at the observatory. This implies that the observing database will not be available until all services are reestablshed. This may be on Aug 17th at the earliest, depending on what progress will be made over the next few days. Also we hope to resume normal night-time operation from the 18th onwards.